# Hydrodynamic noise in one dimension: projected Kubo formula and its vanishing in integrable models

### v1

@@ 5 Absence of noise in integrable systems
As $n$ can be taken as large as we wish and $T^{\prime}>1$, this implies $\hat{\mathfrak{L}}_{j_{i},j_{i}}=0$, and by Cauchy Schwartz shows (91).
@@ 6 Conclusion
We have developed a hydrodynamic fluctuation theory for linearly degenerate systems in one spatial dimension, up to, including, the diffusive order $1/\ell$ as $\ell\to\infty$ in the ballistic scaling of space and time, ${\rm x}=\ell x,\,\mathrm{t}=\ell t$. The theory is presented as a stochastic PDE in a $1/\ell$ expansion, which is well defined up to the first subleading order. This is based on a nonlinear fluctuating Boltzmann-Gibbs principle, where generic observables are projected onto conserved densities, with the addition of diffusive and noise terms. This nonlinear fluctuating Boltzmann-Gibbs principle is fully justified by heuristic arguments based on the separation of relaxation time-scales between coarse-grained conserved densities and other coarse-grained observables. Relaxation of arbitrary observables to their microcanonical averages is modified by noise via the central limit theorem, by “bare” diffusion via fluctuation-dissipation, and by a non-fluctuating, non-diffusive correction to the microcanonical average, in inverse power of the size of the fluid cell. The noise covariance is shown to be given by a projected Kubo formula, where diffuve-scale effects of long-range correlations [19] are subtracted (projected out, from a Hilbert space perspective [10]). The diffusion is shown to be obtained by Einstein relation from this, and the microcanonical correction is shown to give a point-splitting regularisation [19] that makes the stochastic PDE, with delta-correlations both in noise and initial conditions, well-defined. We show, in two different ways, that the noise vanishes at the diffusive scale in integrable models, proving conjectures made previously [20, 21, 19]. We write down the full hydrodynamic equation for linearly degenerate systems, including effects of long-range correlations [19] and bare diffusion.
@@
It would be interesting to evaluate the first subleading corrections to correlation functions beyond the Euler (large-deviation) scale, and to ballistic large-deviation functions [24, 25]. It would also be interesting to generalise this to higher dimensions (see [26] for the nonlinear projection principles for correlation functions at the leading Euler scale), where long-range correlations do not affect the diffusive scale but may affect higher scales of the hydrodynamic expansion.
It would also be interesting to extend the present theory to higher orders. There, $\mathfrak{L}$ and $\mathfrak{D}$ will need regularisation, which will need to be specified – does point-splitting still work? A natural conjecture is that in integrable systems, currents $j_{i}$ do not show emergent noise at all orders in the hydrodynamic expansion in $1/\ell$,
$\overline{o}(x,t)\ \mbox{in $\langle\cdots\rangle_{\ell}$}\longrightarrow o(x,t):=\underbrace{\mathsf{o}^{\rm reg}(\underline{q}(x,t))}_{\text{microcanonical}}+\,\mathcal{O}(\ell^{-\infty})\ \mbox{in $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$}.$ (103)
This is related to the natural conjecture that semiclassical Bethe particles [43, 44, 45] give a good representation of generalised hydrodynamics at all orders [45, 46]. It would be interesting to prove this within the present formalism.
@@
Acknowledgments. I am grateful to Friedrich Hübner and Tomohiro Sasamoto for discussions, and especially Takato Yoshimura for exchanging ideas while working on the simultaneous paper [23] where related results are obtained. I thank the Institute of Science Tokyo for hospitaly while part of this work was done. This work was supported by EPSRC under grants EP/W010194/1, and EP/Z534304/1 (ERC advanced grant scheme).

### v2

@@ 5 Absence of noise in integrable systems
As $n$ and $T^{\prime}>0$ can be taken as large as we wish, this implies $\hat{\mathfrak{L}}_{j_{i},j_{i}}=0$, and by Cauchy Schwartz shows (94).
We note that the presence of a single higher time flow, $n=1$, is sufficient for this argument, as we can take $T^{\prime}>1$ as large as we wish (but according to integrability theory, a single higher-flow implies the presence of infinitely-many higher flows).
Now let us consider the observable $j_{i}]$. Because the higher flows commute, this is a good current observable, for a conserved density that is likewise averaged over higher time flows:
$\partial_{\rm x}[q_{i}]+\partial_{\mathrm{t}}[j_{i}]=0.$ (106)
The above argument shows that as we make $n$ larger for any fixed $T^{\prime}>1$, the emergent noise’s variance becomes smaller. By the Bienaymé-Chebyshev theorem, this implies that all higher cumulants are likewise made smaller. Therefore, we find that the following particular choice of densities and currents,
$q_{i}]_{\infty}=\lim_{n\to\infty}[q_{i}],\quad[j_{i}]_{\infty}=\lim_{n\to\infty}[j_{i}]$ (107)
are good choices of gauge that guarantee that the associated current in the fluctuating hydrodynamic theory has vanishing noise at all orders of the hydrodynamic expansion, and we obtain (91).
Remark 5.1.
Note how this is related to the natural conjecture that semiclassical Bethe particles [66, 67, 68] give a good representation of generalised hydrodynamics at all orders [68, 69], although still missing is the connection between the above choice of gauge and the choice that associates currents to these particles’ flows.
6 Hydrodynamic equation and diffusive two-point functions
Sec. 4.2 and 4.3 illustrate some of the techniques involved in calculating using the nonlinear fluctuating Boltzmann-Gibbs principle (21), in particular combining the point-splitting regularisation with the cumulant expansion. In order to further illustrate this “calculus” and to show the general consistency of our framework, in this section we provide two simple applications. We obtain the hydrodynamic equation up to the diffusive correction, first for one-point function in non-equilibrium states emanating from (12), and second for two-point functions of conserved densities in stationary states.
For one-point function, the result is highly anomalous, taking the form of a coupled system of equations as in [25] for integrable systems, but generalised to the presence of bare diffusion. The calculation is a simple generalisation of that shown in the main text of [25].
For two-point functions, the result is a normal diffusion equation, with the full diffusion matrix related by Einstein relation to the full, un-projected Onsager matrix. This is usually assumed to be valid and follows from sum-rule manipulations, see e.g. [5, 41]. Here we derive it from our fluctuating hydrodynamic theory, in which context it is somewhat non-trivial, as the effects of long-range correlations to the diffusive scale – coming from the regularised microcanonical part of (21) – and of bare diffusion must combine just in the right way to give the diffusion matrix related to the full Onsager matrix. This was also derived in [25] for integrable systems, although in this case there is no bare diffusion. It was derived by a different technique, using a special linear response theory on the hydrodynamic equation for one-point functions in non-equilibriun states. There, the full Onsager matrix – without bare diffusion – arises because the at infinitesimal macroscopic times, the long-range correlations exactly reproduce the projection onto quadratic charges, thus, for integrable systems the full Onsager matrix (as we showed in Sec. 5). We instead perform a direct calculation using the nonlinear fluctuating Boltzmann-Gibbs principle, which involves 3-point functions in stationary states recently evaluated [43].
Thus we show that the present framework gives a consistent theoretical underpinning for both the special emergent hyrodynamic equation out of equilibrium, and normal diffusion at the linearised level.
6.1 The hydrodynamic equations at the diffusive scale
Using our fluctuating hydrodynamic theory based on (21), we first write down the diffusive-scale hydrodynamic equation, taking into account the results Points I-III. For definiteness the initial state is the long-wavelength state (12), but more general states would lead to the same equation.
The hydrodynamic equation is obtained from the conservation laws:
Inserting (21) for the case $o=j_{i}$ in this, using $\langle\!\langle\xi_{o}(x,t)\rangle\!\rangle_{\ell}=0$ and the fact that for the diffusive order $\ell^{-1}$ we may use the leading cumulant-expansion form
$\partial_{t}\mathsf{q}_{i}(x,t)+\partial_{x}\langle\!\langle\mathsf{j}_{i}^{\rm reg}(\underline{q}(x,t))\rangle\!\rangle_{\ell}=\frac{1}{2\ell}\sum_{k}\partial_{x}\Big{(}\hat{\mathfrak{D}}_{i}^{~k}(\underline{\mathsf{q}}(x,t))\partial_{x}\mathsf{q}_{k}(x,t)\Big{)}$ (110)
$\partial_{t}\mathsf{q}_{i}(x,t)+\sum_{k}\mathsf{A}_{i}^{~k}(x,t)\partial_{x}\mathsf{q}_{k}(x,t)+\frac{1}{4\ell}\sum_{I,J,\pm}\partial_{x}\Big{(}\langle j_{i}^{-},Q_{I},Q_{J}\rangle^{\rm c}_{\underline{\beta}(x,t)}E_{IJ}(x\pm 0^{+},x\mp 0^{+};t)\Big{)}$
$\hskip-156.49014pt=\frac{1}{2\ell}\sum_{k}\partial_{x}\Big{(}\hat{\mathfrak{D}}_{i}^{~k}(\underline{\mathsf{q}}(x,t))\,\partial_{x}\mathsf{q}_{k}(x,t)\Big{)}$
with
$\partial_{t}E_{IJ}(x,y;t)+\partial_{x}\big{(}\mathsf{v}_{I}(x,t)E_{IJ}(x,y;t)\big{)}+\partial_{y}\big{(}\mathsf{v}_{J}(y,t)E_{IJ}(x,y;t)\big{)}$
$\hskip 170.71652pt=\delta(x-y)\sum_{k}\langle j_{k}^{-},Q_{I},Q_{J}\rangle^{\rm c}_{\underline{\beta}(x,t)}\partial_{x}\beta^{k}(x,t).$
Here $\beta^{i}(x,t):=\beta^{i}(\underline{\mathsf{q}}(x,t))$ (see Eq. (11)) and likewise for the hydrodynamic velocities $\mathsf{v}_{I}(x,t)$. In the state (12), the initial conditions are
$\beta^{i}(x,0)=\beta^{i}(x),\quad E_{rs}(x,y;0)=0$ (113)
but these equations are more general, and valid also for initial states that admit non-trivial long-range correlations $E_{rs}(x,y;0)\neq 0$. Eqs. (LABEL:hydro1), (LABEL:hydro2) generalise [25, Eqs 9, 10], which was for integrable models, to the inclusion of the bare diffusion generically present away from integrability.
For integrable models, Eqs. (LABEL:hydro1) and (LABEL:hydro2) are immediate to write in terms of quasi-particle rapidities, using that, in normal modes $I\in{\mathbb{R}}$ (which are of rapidity type), $\mathsf{A}_{I}^{~J}=\delta(I-J)v^{\rm eff}(I)$, and the three-point coupling in normal modes (143) has the expression displayed in [43, Eq 103]. See [69] where it is written in even more generality, including external force terms [70].
6.2 Normal diffusion for stationary-state two-point functions
We restrict to stationary states
$\langle\cdots\rangle=\langle\cdots\rangle_{\underline{\beta}}$ (114)
and all constitutive functions are evaluated within this state (we keep their state dependence implicit).
We will show that the diffusive equation for two-point functions of conserved densities takes a diffusive form:
$\partial_{t}\langle\overline{q_{i}}(x,t),\overline{q_{j}}(0,0)\rangle^{\rm c}+\sum_{k}\mathsf{A}_{i}^{~k}\partial_{x}\langle\overline{q_{k}}(x,t),\overline{q_{j}}(0,0)\rangle^{\rm c}=\frac{1}{2\ell}\sum_{k}\mathfrak{D}_{i}^{~k}\partial_{x}^{2}\langle\overline{q_{k}}(x,t),\overline{q_{j}}(0,0)\rangle^{\rm c},\quad t>0.$ (115)
This is in principle valid for fluid-cell means as written here, so that possible oscillatory behaviours are averaged out, but in practice often the fluid-cell mean is not required. Importantly, as mentioned, the diffusion matrix involved is related by the Einstein relation to the full Onsager matrix,
$\mathfrak{L}_{il}=\sum_{k}\mathfrak{D}_{i}^{~k}\mathsf{C}_{kl}.$ (116)
Here and below we use the simplified notation $\mathfrak{L}_{ik}=\mathfrak{L}_{j_{i},j_{k}}$ paralleling (27). Eq. (116) is the same relation as Eq. (47) for the bare diffusion matrix in terms of the projected Kubo formula.
That the full diffusion matrix be involved in the two-point function is non-trivial: the bare diffusion term easily gives a contribution to the diffusion term in (115), however the long-range correlations, coming from the regularised microcanonical average, also give a contribution, and these sum up exactly to the full diffusion matrix. This is what we will show now. The proof is based on a result for the exact three-point function of conserved densities at the Euler scale that we obtained recently by projection techniques, see [43, Sec 5.2].
Using the conservation laws again, passing to the fluctuating theory, and using normal modes and the fact that two-point functions of conserved densities at the Euler scale are diagonal in this basis, Eq. (152), we must show that
$\langle\!\langle j_{I}(x,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}=\delta_{IJ}\mathsf{v}_{I}\langle\!\langle q_{I}(x,t),q_{I}(0,0)\rangle\!\rangle^{\rm c}-\frac{1}{2\ell}\mathfrak{L}_{IJ}\partial_{x}\langle\!\langle q_{J}(x,t),q_{J}(0,0)\rangle\!\rangle^{\rm c},\quad t>0.$ (117)
As the second term on the right-hand side has a factor $1/\ell$, it can be evaluated at the Euler scale, so we must show
$\langle\!\langle j_{I}(x,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}=\delta_{IJ}\mathsf{v}_{I}\langle\!\langle q_{I}(x,t),q_{I}(0,0)\rangle\!\rangle^{\rm c}-\frac{1}{2\ell^{2}}\mathfrak{L}_{IJ}\delta^{\prime}(x-\mathsf{v}_{J}t),\quad t>0.$ (118)
Proof.
We use (21) for $o=j_{I}$. Using (32), as the noise $\xi_{I}$ does not correlate with the fluctuation of $q_{J}(0,0)$, induced by the initial state, it does not contribute:
$\langle\!\langle j_{I}(x,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}\stackrel{{\scriptstyle\text{noise}}}{{=}}0.$ (119)
The diffusive term can be evaluated by the cumulant expansion, at leading order giving simply
$\langle\!\langle j_{I}(x,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}\stackrel{{\scriptstyle\text{diffusive}}}{{=}}-\frac{1}{2\ell}\hat{\mathfrak{L}}_{IJ}\partial_{x}\langle\!\langle q_{J}(x,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}=-\frac{1}{2\ell^{2}}\hat{\mathfrak{L}}_{IJ}\delta(x-\mathsf{v}_{J}t),\quad t>0.$ (120)
This is the bare contribution to the diffusive term of the two-point function.
The most important part is that coming from the microcanonical average, with point-splitting regularisation. Using (4.2), this is
$\langle\!\langle\mathsf{j}_{I}^{\rm reg}(\underline{q}(x,t))q_{J}(0,0)\rangle\!\rangle^{\rm c}$
$=$ $\delta_{IJ}\mathsf{v}_{I}\langle\!\langle q_{I}(x,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}+\frac{1}{2}\sum_{KL}\mathsf{A}_{I}^{~KL}\frac{1}{2}\sum_{\epsilon=\pm 0^{+}}\langle\!\langle q_{K}(x+\epsilon,t),q_{L}(x-\epsilon,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}.$
We must now evaluate the three-point function. For this purpose, it is simpler to bring the space-time variables $x,t$ onto $q_{J}$ by translation invariance and PT symmetry:
$\langle\!\langle q_{K}(x+\epsilon,t),q_{L}(x-\epsilon,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}=\langle\!\langle q_{K}(-\epsilon,0),q_{L}(+\epsilon,0),q_{J}(x,t)\rangle\!\rangle^{\rm c}.$ (122)
We then use the formula [43, Eqs 98, 99]. This has two types of contribution:
$\langle\!\langle q_{K}(-\epsilon,0),q_{L}(\epsilon,0),q_{J}(x,t)\rangle\!\rangle^{\rm c}=\ell^{-2}(C_{\rm linear}-C_{\rm nonlinear}).$ (123)
The first is a direct, linear propagation. This contribution vanishes because of the point-splitting:
$C_{\rm linear}=\langle Q_{K},Q_{L},q_{J}\rangle^{\rm c}\delta(x+\epsilon-\mathsf{v}_{J}t)\delta(x-\epsilon-\mathsf{v}_{J}t)=0.$ (124)
The second is an indirect, nonlinear propagation, coming from the nonlinear part of the 3-point projection. Different cases in principle must be treated, depending if some of the velocities $\mathsf{v}_{J},\,\mathsf{v}_{K},\,\mathsf{v}_{L}$ are equal to each other. However, the expression in the case where only two velocities are equal to each other, can be obtained by taking the formal limit of equal velocities; while that where all velocities are equal to each other does not contribute because of the vanishing 3-point coupling (161). Hence, it is sufficient to perform the full calculation from the expression for unequal velocities. Because of our choice of values of times, only the nonlinear propagation associated to $q_{J}$ is required, giving:
$C_{\rm nonlinear}=\mathsf{A}_{J}^{~KL}\delta^{\prime}(v^{\prime}u-vu^{\prime})\big{(}|v^{\prime}|(s(u^{\prime}+v^{\prime}t)-s(u^{\prime}))-|v|(s(u+vt)-s(u))\big{)}=\mathsf{A}_{J}^{~KL}c$ (125)
where $s(a)=\frac{1}{2}\operatorname{sgn}(a)$ and
$u=x-\mathsf{v}_{J}t+\epsilon,\ u^{\prime}=x-\mathsf{v}_{J}t-\epsilon,\ v=\mathsf{v}_{JK},\ v^{\prime}=\mathsf{v}_{JL}$ (126)
with the assumption $\mathsf{v}_{J}\neq\mathsf{v}_{K},\,\mathsf{v}_{J}\neq\mathsf{v}_{L},\,\mathsf{v}_{K}\neq\mathsf{v}_{L}$. Thus we must evaluate:
$c$ $=$ $\delta^{\prime}\big{(}\mathsf{v}_{KL}(x-\mathsf{v}_{J}t)+\epsilon(2\mathsf{v}_{J}-\mathsf{v}_{K}-\mathsf{v}_{L})\big{)}\times$
$\times\,\Big{(}|\mathsf{v}_{JL}|(s(x-\mathsf{v}_{L}t-\epsilon)-s(x-\mathsf{v}_{J}t-\epsilon))-|\mathsf{v}_{JK}|(s(x-\mathsf{v}_{K}t+\epsilon)-s(x-\mathsf{v}_{J}t+\epsilon))\Big{)}$
We evaluate the term with $s(x-\mathsf{v}_{J}t-\epsilon)$ as
[…]
$\stackrel{{\scriptstyle|\epsilon|\to 0}}{{=}}$ $\frac{|\mathsf{v}_{JL}|}{\mathsf{v}_{KL}|\mathsf{v}_{KL}|}\Big{(}\delta^{\prime}(x-\mathsf{v}_{J}t)s\Big{(}\frac{\mathsf{v}_{JL}}{\mathsf{v}_{KL}}\Big{)}\operatorname{sgn}(\epsilon)+\delta(x-\mathsf{v}_{J}t)\delta\Big{(}\frac{\mathsf{v}_{JL}}{\mathsf{v}_{KL}}\epsilon\Big{)}\Big{)}$
$=$ $\frac{|\mathsf{v}_{JL}|}{\mathsf{v}_{KL}|\mathsf{v}_{KL}|}\delta^{\prime}(x-\mathsf{v}_{J}t)s\Big{(}\frac{\mathsf{v}_{JL}}{\mathsf{v}_{KL}}\Big{)}\operatorname{sgn}(\epsilon)$
and under the sum $\sum_{\epsilon=\pm 0^{+}}$ in (6.2) the result vanishes. Likewise for $s(x-\mathsf{v}_{J}t+\epsilon)$,
$\frac{|\mathsf{v}_{JK}|}{\mathsf{v}_{KL}|\mathsf{v}_{KL}|}\delta^{\prime}\Big{(}x-\mathsf{v}_{J}t+\epsilon\frac{2\mathsf{v}_{J}-\mathsf{v}_{K}-\mathsf{v}_{L}}{\mathsf{v}_{KL}}\Big{)}s(x-\mathsf{v}_{J}t+\epsilon)\stackrel{{\scriptstyle|\epsilon|\to 0}}{{=}}-\frac{|\mathsf{v}_{JK}|}{\mathsf{v}_{KL}|\mathsf{v}_{KL}|}\delta^{\prime}(x-\mathsf{v}_{J}t)s\Big{(}\frac{\mathsf{v}_{JK}}{\mathsf{v}_{KL}}\Big{)}\operatorname{sgn}(\epsilon)$ (129)
which again vanishes under $\sum_{\epsilon=\pm 0^{+}}$. For the term with $s(x-\mathsf{v}_{L}t-\epsilon)$ we get
$\frac{|\mathsf{v}_{JL}|}{\mathsf{v}_{KL}|\mathsf{v}_{KL}|}\delta^{\prime}\Big{(}x-\mathsf{v}_{J}t+\epsilon\frac{2\mathsf{v}_{J}-\mathsf{v}_{K}-\mathsf{v}_{L}}{\mathsf{v}_{KL}}\Big{)}s(x-\mathsf{v}_{L}t-\epsilon)$ (130)
$\stackrel{{\scriptstyle|\epsilon|\to 0}}{{=}}$ $\frac{|\mathsf{v}_{JL}|}{\mathsf{v}_{KL}|\mathsf{v}_{KL}|}\big{(}\delta^{\prime}(x-\mathsf{v}_{J}t)s(\mathsf{v}_{JL}t)-\delta(x-\mathsf{v}_{J}t)\delta(\mathsf{v}_{JL}t)\big{)}$
$\stackrel{{\scriptstyle t>0}}{{=}}$ $\frac{\mathsf{v}_{JL}}{2\mathsf{v}_{KL}|\mathsf{v}_{KL}|}\delta^{\prime}(x-\mathsf{v}_{J}t)$
and likewise for the term with $s(x-\mathsf{v}_{K}t+\epsilon)$,
$\frac{|\mathsf{v}_{JK}|}{\mathsf{v}_{KL}|\mathsf{v}_{KL}|}\delta^{\prime}\Big{(}x-\mathsf{v}_{J}t+\epsilon\frac{2\mathsf{v}_{J}-\mathsf{v}_{K}-\mathsf{v}_{L}}{\mathsf{v}_{KL}}\Big{)}s(x-\mathsf{v}_{K}t+\epsilon)\stackrel{{\scriptstyle|\epsilon|\to 0,\,t>0}}{{=}}-\frac{\mathsf{v}_{JK}}{2\mathsf{v}_{KL}|\mathsf{v}_{KL}|}\delta^{\prime}(x-\mathsf{v}_{J}t).$ (131)
Combining,
$\frac{1}{2}\sum_{\epsilon=\pm 0^{+}}C_{\rm nonlinear}=\mathsf{A}_{J}^{~KL}\frac{1}{2|\mathsf{v}_{KL}|}\delta^{\prime}(x-\mathsf{v}_{J}t)$ (132)
and therefore
$\frac{1}{2}\sum_{KL}\mathsf{A}_{I}^{~KL}\frac{1}{2}\sum_{\epsilon=\pm 0^{+}}\langle\!\langle q_{K}(x+\epsilon,t),q_{L}(x-\epsilon,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}\stackrel{{\scriptstyle t>0}}{{=}}-\sum_{KL}\frac{\mathsf{A}_{I}^{~KL}\mathsf{A}_{J}^{~KL}}{4|\mathsf{v}_{KL}|\ell^{2}}\delta^{\prime}(x-\mathsf{v}_{J}t).$ (133)
With (6.2), (120) and (133), we therefore find
$\langle\!\langle j_{I}(x,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}=\delta_{IJ}\mathsf{v}_{I}\langle\!\langle q_{I}(x,t),q_{J}(0,0)\rangle\!\rangle^{\rm c}-\frac{1}{2\ell^{2}}\Big{(}\hat{\mathfrak{L}}_{IJ}+\sum_{KL}\frac{\mathsf{A}_{I}^{~KL}\mathsf{A}_{J}^{~KL}}{2|\mathsf{v}_{KL}|}\Big{)}\delta^{\prime}(x-\mathsf{v}_{J}t).$ (134)
Thus from (44) and (143), we have shown that (118) holds.
@@ 7 Conclusion
We have developed a hydrodynamic fluctuation theory for linearly degenerate systems in one spatial dimension, up to, including, the diffusive order $1/\ell$ as $\ell\to\infty$ in the ballistic scaling of space and time, ${\rm x}=\ell x,\,\mathrm{t}=\ell t$. The theory is presented as a stochastic PDE in a $1/\ell$ expansion, which is well defined up to the first subleading order. This is based on a nonlinear fluctuating Boltzmann-Gibbs principle, where generic observables are projected onto conserved densities, with the addition of diffusive and noise terms. This nonlinear fluctuating Boltzmann-Gibbs principle is justified by heuristic arguments that we have disccused carefully, based on the separation of relaxation time-scales between coarse-grained conserved densities and other coarse-grained observables. Relaxation of arbitrary observables to their microcanonical averages is modified by noise via the central limit theorem, by “bare” diffusion via because of relaxation time delay (fluctuation-dissipation), and by a non-fluctuating, non-diffusive correction to the microcanonical average, representing the correction to the microcanonical average due to the finite size of the region (the fluid-cell) on which it is taken. The noise covariance is shown to be given by a projected Kubo formula, where diffusive-scale effects of long-range correlations [25] are subtracted (projected out, from a Hilbert space perspective [24]). The diffusion is shown to be obtained by Einstein relation from this, and the microcanonical correction is shown to give a point-splitting regularisation [25] that makes the stochastic PDE, with delta-correlations both in noise and initial conditions, well-defined. We show, in two different ways, that the noise vanishes at the diffusive scale in integrable models, proving conjectures made previously [36, 37, 25]. We show that, under an appropriate choice of currents, their noise in fact vanishes at all orders in the hydrodynamic expansion. We write down the full hydrodynamic equation for linearly degenerate systems, including effects of long-range correlations [25] and bare diffusion. We also derived, within our fluctuating hydrodynamic formalism, the diffusive hydrodynamic equation for dynamical correlation functions in stationary states, where we showed that diffusion is associated to the full Onsager matrix instead of the projected one, as it should be.
In the heuristic arguments we have provided to explain the form of the nonlinear Boltzmann-Gibbs principle in Sec. 3.2, noise emerges under the usual assumption of molecular chaos, here expressed as the assumption that interaction events are essentially independent and non-correlated when separated enough in space and time (see Fig. 2). Clearly, as we show that noise in fact vanishes for currents in integrable models, there may in fact be correlations that reduce the noise even for certain observables which are not conserved densities. In Hamiltonian systems, one of the leading open problems in the hydrodynamic theory remains to more accurately understand the emergence of Gaussian white noise, perhaps by expanding on the heuristic arguments we have expressed.
In order to make the absence of noise for currents in integrable models more explicit, and to put the heuristic arguments of Sec. 3.2 on a more solid basis, one should investigate numerically fluctuations of observables on fluid cells [^6] As we mentioned, shortly after the first version of this paper was posted, paper [46] performed this in the hard rods model, confirming our results..
It has recently been shown with mathematical rigour that in (generalised) hard rods models, a certain fluctuating hydrodynamic equation emerges for linear fluctuating fields, where noise appears, albeit with strong correlations in space [71]. This noise is not of the type that appears in (21), which, as we have shown, vanishes in integrable models. Instead, we believe this space-correlated noise is a “remnant”, at the linear level, of long-range correlations that are known to occur from the present nonlinear version of the fluctuating hydrodynamics. More investigation on this aspect would be appropriate.
@@
It would be interesting to evaluate explicitly, using the present theory, the first subleading corrections to correlation functions and to large-deviation functions beyond the Euler scale [7, 8, 43]. It would also be interesting to generalise this to higher dimensions (see [43] for the nonlinear projection principles for correlation functions at the leading Euler scale), where long-range correlations do not affect the diffusive scale but may affect higher scales of the hydrodynamic expansion.
@@
Acknowledgments. I am grateful to Friedrich Hübner, Tomohiro Sasamoto and Mayank Sharma for discussions, and especially Takato Yoshimura for exchanging ideas while working on the simultaneous paper [39] where related results are obtained. I thank the Institute of Science Tokyo for hospitaly, and the workshop “Hydrodynamics of low-dimensional interacting systems: Advances, challenges, and future directions", YITP-T-25-03, at the Yukawa Institute for Theoretical Physics, where part of this work was done. This work was supported by EPSRC under grants EP/W010194/1, and EP/Z534304/1 (ERC advanced grant scheme).